# GaussRender: Learning 3D Occupancy with Gaussian Rendering

### v2

@@ 4.3 Finer-grained multi-view metric analysis
[…]
TPVFormer [15] 58.16 28.48
SurroundOcc- w/ GaussRender 59.20 +1.04 28.73 +0.25
nuSc [54] SurroundOcc [54] 58.60 28.26
w/ GaussRender 60.55 +1.95 28.64 +0.38
TPVFormer [15] 52.95 29.72
Occ3D- w/ GaussRender 54.35 +1.40 30.26 +0.54
nusc [50] SurroundOcc [54] 53.52 28.98
w/ GaussRender 55.65 +2.13 30.50 +1.52
[…]
@@
BeV metrics.
[…]
Our analysis is presented in Tab. 3. We observe that the use of GaussRender enhances both metrics simultaneously, with systematic gains associated with different combinations of datasets and models. Both TPVFormer and SurroundOcc show significant improvements across all datasets and evaluations. This evaluation highlights that the use of GaussRender not only improves 3D occupancy predictions but also enhances consistency with BeV and 2D sensor observations.
@@ 4.4 Ablation studies
SurroundOcc-nuSc [54] Occ3d-nuSc [50]
Visible voxels only
Camera positioning IoU ($\uparrow$) mIoU ($\uparrow$) IoU ($\uparrow$) mIoU ($\uparrow$)
Sensor 8.7 0.8 41.1 16.1
Elevated 24.9 7.1 42.0 11.7
Random 24.5 6.9 38.4 12.0
Stereo 24.8 8.5 41.1 15.8
Elevated + random 25.4 8.5 40.8 12.9
Table 4: Impact of different camera positioning strategies on 3D semantic occupancy performances. The architecture used is TPVFormer [15]. Models are trained and evaluated on 20% of the SurroundOcc-nuScenes and Occ3d-nuScenes datasets [54, 50] using only $L_{\text{2D}}$, without $L_{\text{3D}}$, but are evaluated with classical 3D IoU and mIoU.
@@ 4.4.1 Impact of supervising with virtual viewpoints
A key advantage of GaussRender is that 3D occupancy can be rendered from arbitrary viewpoints. To explore the impact of different virtual camera placements, we evaluate several positioning strategies described below:
  • Sensor Strategy: Cameras are placed at the original positions and orientations from the dataset.
  • Random Strategy: Cameras are randomly positioned within the scene by applying a pitch and yaw variation of $\pm$10° and a forward-backward shift from the half of the maximum range.
  • Elevated Strategy: Cameras are lifted by 2 meters while tilted 20° downward. This provides a wider field of view and reduces voxel occlusion.
  • Elevated + Random Strategy: Cameras are elevated by 2 meters, tilted 20° downward, and randomly moved in the scene around the vehicle to half of the maximum range. This allows capturing voxels that are not visible from static sensor positions.
  • Stereo Strategy: We select one sensor camera and position another one close to the sensor to mimic a stereo behavior.
Tab. 4 quantifies the impact of these strategies on the final 3D predictions. The choice of camera placement should be tailored to the specific task. If the goal is to predict only visible voxels, as in Occ3D-nuScenes [50], using sensor-aligned viewpoints yields the best results. However, if the objective is to also infer occluded regions, as in SurroundOcc-nuScenes [54], using sensor-aligned viewpoints leads to low results but placing virtual cameras to maximize the visibility of these areas provides better supervision. The random strategy further highlights the importance of careful camera positioning: placing cameras arbitrarily often leads to poor 3D metrics, as they mainly observe empty space.
@@ Appendix C Computational cost
As reported in Tab. 6, training with GaussRender incurs a modest increase in memory and computation time ( 10%), even with high-resolution renderings. This overhead can be further reduced by pre-selecting camera locations, allowing annotation renderings (the two ground-truth renderings) to be pre-processed in advance. Additionally, lower rendering resolutions can be used if needed.
@@
Appendix D Ablations
D.1 Gaussian scaling
@@ D.3 Cameras
Appendix E Scores detailed per class
The following tables give the detailed IoU and mIoU scores of the models studied for each dataset. Tab. 8 concerns the Occ3D-nuScenes dataset [50], Tab. 9 the SSCBench-KITTI360 dataset [31] and Tab. 7 the SurroundOcc-nuScenes dataset [54].
Appendix F Qualitative results

### v3

@@ 4.3 Finer-grained multi-view metric analysis
Models RayIoU RayIoU1m, 2m, 4m
RenderOcc [43] 19.5 13.4, 19.6, 25.5
BEVDet-Occ [13] 29.6 23.6, 30.0, 35.1
BEVFormer [32] 32.4 26.1, 32.9, 38.0
BEVDet-Occ-Long [13] 32.6 26.6, 33.1, 38.2
SparseOcc[36] (8f) 34.0 28.0, 34.7, 39.4
FB-Occ [33] 33.5 26.7, 34.1, 39.7
SparseOcc [36] (16f) 36.1 30.2, 36.8, 41.2
TPVFormer [15] 37.2 31.5, 38.1, 41.9
w/ GaussRender 38.3 +1.1 32.3, 39.3, 43.4
SurroundOcc [54] 35.5 29.9, 36.3, 40.1
w/ GaussRender 37.5 +2.0 31.4, 38.5, 42.6
Table 4: Impact of GaussRender on RayIoU ($\uparrow$) metrics on the Occ3D-nuScenes validation dataset. Best results are in bold. Previous results are reported from [36]. RayIoU1m, 2m, 4m refers to the RayIoU with a depth tolerance of 1, 2, or 4 meters.
@@ 4.4 Ablation studies
Cameras Sensor Elevated Elevated + Around. Fully Rand. Dynamic
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
2D+3D 25.9 26.0 26.3 25.4 25.8
2D-only 16.1 11.7 12.9 6.9 4.8
Table 5: Comparison of mIoU with different camera sampling strategies trained and evaluated on 20% of Occ3D. We evaluate five sampling strategies under 2D-only and 2D+3D supervision. Images illustrate the impact of camera positioning, black dots represent camera centers while red dots represent visible points casted in their frustum.
@@ 4.4.1 Impact of supervising with virtual viewpoints
A key advantage of GaussRender is its ability to render 3D occupancy from arbitrary viewpoints, offering a flexible way to supervise 3D occupancy. To investigate the influence of virtual camera configurations, we evaluate five different camera placement strategies: Sensor, Elevated, Elevated + Around, Fully Random, Dynamic. More details in Appendix E.4. Quantitatively Tab. 5 quantifies the effect of each strategy under both 2D-only and 2D+3D supervision using mIoU. To do this we train a TPVFormer model with GaussRender during 20% of the Occ3d-nuSc training set. Several conclusions can be drawn. Under 2D-only supervision, the Sensor strategy leads by a large margin (16.1 mIoU), suggesting that when only 2D supervision is available, aligning virtual views with actual camera positions maximizes consistency and learning efficiency. Both Fully Random and Dynamic strategies suffer heavily, with performance dropping to 6.9 and 4.8 mIoU, respectively. These results highlight the risk of unstructured camera placement: random viewpoints often observe unoccupied or irrelevant parts of the scene, weakening the supervision signal while dynamic positioning may introduce instability and bad signals during training leading to poor results. Under 2D+3D supervision, we clearly see that what matters is to have complementary viewpoints since the Sensor strategy is beaten by Elevated + Around achieving the best result (26.3 mIoU). This indicates that providing diverse and informative viewpoints enhances learning, even when 3D supervision is available.
Overall, these findings emphasize that virtual camera placement is important. If we consider a 2D only training, positioning cameras at sensor location ensures consistency with sensors during training, while when training 2D+3D, we need additional viewpoints to enhance supervision.
@@ Appendix C Computational cost
As reported in Tab. 7, training with GaussRender incurs a modest increase in memory and computation time ( 10%), while using high-resolution renderings. This overhead can be further reduced by pre-selecting camera locations, allowing annotation renderings (the two ground-truth renderings) to be pre-processed in advance. Additionally, lower rendering resolutions can be used if needed.
@@
Appendix D BeV metrics.
[…]
TPVFormer [15] 58.16 28.48
SurroundOcc- w/ GaussRender 59.20 +1.04 28.73 +0.25
nuSc [54] SurroundOcc [54] 58.60 28.26
w/ GaussRender 60.55 +1.95 28.64 +0.38
TPVFormer [15] 52.95 29.72
Occ3D- w/ GaussRender 54.35 +1.40 30.26 +0.54
nusc [50] SurroundOcc [54] 53.52 28.98
w/ GaussRender 55.65 +2.13 30.50 +1.52
[…]
Our analysis is presented in Tab. 8. We observe that the use of GaussRender enhances both metrics simultaneously, with systematic gains associated with different combinations of datasets and models. Both TPVFormer and SurroundOcc show significant improvements across all datasets: Occ3d-nuScenes and SurroundOcc-nuScenes and evaluations. This evaluation highlights that the use of GaussRender not only improves 3D occupancy predictions but also enhances consistency with BeV and sensor observations.
Appendix E Ablations
E.1 Gaussian scaling
@@ E.3 Cameras
E.4 Camera strategies
  • Sensor Strategy: Cameras are placed at the original sensor locations and orientations as provided in the dataset.
  • Elevated Strategy: Each camera is lifted and tilted downward. This modification increases the vertical field of view, providing a top-down perspective that reduces self-occlusion and captures a broader context of the scene.
  • Elevated + Around Strategy: This strategy combines elevation and downward tilt with additional random displacements around the ego vehicle—up to half the maximum scene range. It allows observing the scene from novel angles while maintaining a consistent overhead viewpoint, improving the visibility of occluded voxels.
  • Fully Random Strategy: Cameras are randomly placed throughout the scene, applying pitch and yaw perturbations and varying distances from the ego-vehicle. While this increases the diversity of viewpoints, it also introduces inconsistency and often places cameras in less informative positions (e.g., viewing empty space).
  • Dynamic Strategy: Cameras are elevated and positioned at random distances along a circular ring. However, each camera is oriented to look at the element with the highest 3D cross-entropy. In other words, the camera focuses on a region where it made the largest prediction error. It appears that its supervision signal does not help much the training.
@@ E.4 Camera strategies
Appendix F Scores detailed per class
The following tables give the detailed IoU and mIoU scores of the models studied for each dataset. Tab. 10 concerns the Occ3D-nuScenes dataset [50], Tab. 11 the SSCBench-KITTI360 dataset [31] and Tab. 9 the SurroundOcc-nuScenes dataset [54]. The variations by class appear to be due to learning variance, which is why it makes more sense to look at the overall IoU and RayIoU metrics, rather than looking for intrinsic reasons.
Appendix G Qualitative results